OpenStudy (anonymous):

My work asks me to identify any extraneous solutions, what exactly am I looking for? Like an alternative answer? (ie: x = 3 OR x = -1)

OpenStudy (anonymous):

Sometimes, for example, a quadratic, produces 2 solutions but one of them is clearly inappropriate (perhaps because it is negative).

OpenStudy (anonymous):

So what about when it's something like -4√(x + 9) = 20

OpenStudy (anonymous):

Should I say something like √(x + 9) = 5 OR √(x + 9) + 5 = 0

OpenStudy (anonymous):

The only way that your equation can have solutions is if sqrt(x+9) = -5 which is impossible, so no solutions exist.

OpenStudy (anonymous):

I see, thanks.

OpenStudy (anonymous):

ur welcome.